If A is the sum of the integers from 1 to 50 and B is the sum of the integers from 51 to 100, what is the value of B – A ?
(A) 2720 (B) 2340 (C) 2500 (D) 1990 (E) 2800

Answer by stanbon(75887)

If A is the sum of the integers from 1 to 50 and B is the sum of the integers from 51 to 100, what is the value of B – A ?
---
A = (50)(1+50)/2 = 25*51
---
B = 50(51+100)/2 = 25(151)
---
B - A = 25(151)-25(51) = 25*100 = 2500
---------------------------------------------
Cheers,
Stan H.
------------
(A) 2720 (B) 2340 (C) 2500 (D) 1990 (E) 2800